# Learning Encodings by Maximizing State Distinguishability: Variational Quantum Error Correction

### v1

@@ F.3 Results on IQM Devices
Figure 24: Reconstructed distinguishability loss for the $3$ patches of the experiment with the [[$4,1$]] VarQEC code on the IQM Garnet device. The qubits have been idled for $1.0\mu\text{s}$ for this experiment. We consider two different physical qubit to logical patch alignment strategies, in particular (a) “star” and (b) “square” alignment. For each patch, the first four bars indicate the unencoded baseline losses for the respective physical qubits, while the fifth one denotes the loss for the entire encoded patch. The losses were estimated with mitigated state tomography using $1000$ shots per circuit execution, and the error bars denote 5th and 95th percentile ranges estimated by cluster re-sampling.
@@
The denser grid connectivity of the IQM device allowed us to experiment with different ansatz topologies, specifically the “star” and “square” layouts. However, the results do not show a significant difference between the two layouts, which we expect to change for larger codes, as analyzed in Section E.4. The increased span of the error bars can be attributed to the reduction of the shot count by a factor of $10$. We observed that in both layouts, qubit $4$ was defective during our experiments, leading to outlier results in patch I. In patches II and III, we achieved a clear net gain in performance, as the logical distinguishability loss was significantly lower than that of the best physical qubit. The smaller variance in qubit quality on the IQM device facilitated better evaluation of our code’s functionality, given that our training assumed uniform noise levels across qubits.

### v2

@@ F.3 Results on IQM Devices
Figure 24: Reconstructed distinguishability loss for the 33 patches of the experiment with the $((4,2))$ VarQEC code on the IQM Garnet device. The qubits have been idled for $1.0\mu\text{s}$ for this experiment. We consider two different physical qubit to logical patch alignment strategies, in particular (a) “star” and (b) “square” alignment. For each patch, the first four bars indicate the unencoded baseline losses for the respective physical qubits, while the fifth one denotes the loss for the entire encoded patch. The losses were estimated with mitigated state tomography using $1000$ shots per circuit execution, and the error bars denote 5th and 95th percentile ranges estimated by cluster re-sampling.
@@
The denser grid connectivity of the IQM device allowed us to experiment with different ansatz topologies, specifically the “star” and “square” layouts. However, the results do not show a significant difference between the two layouts, which we expect to change for larger codes, as analyzed in section˜E.4. The increased span of the error bars can be attributed to the reduction of the shot count by a factor of $10$. We observed that in both layouts, qubit $4$ was defective during our experiments, leading to outlier results in patch I. In patches II and III, we achieved a clear net gain in performance, as the logical distinguishability loss was significantly lower than that of the best physical qubit. The smaller variance in qubit quality on the IQM device facilitated better evaluation of our code’s functionality, given that our training assumed uniform noise levels across qubits.